# Towards Any-Quality Image Segmentation via Generative and Adaptive Latent Space Enhancement

### v1

@@ 5.3.1 Effect of Each Component
As summarized in Tab. 5, we conduct an ablation study to assess the individual contributions of the proposed modules in GleSAM++, namely the Feature Distribution Alignment (FDA), Channel Replication and Expansion (CRE), and Degradation-aware Adaptive Enhancement (DAE). Here, “Gle” denotes the base framework equipped with generative latent space enhancement. To ensure a fair comparison of reconstruction capability, all experiments are performed by fine-tuning only the U-Net backbone. The quantitative results demonstrate that each additional module consistently improves segmentation accuracy across all degradation levels and datasets. Notably, integrating FDA and CRE significantly bridges the gap between the pre-trained latent diffusion model and SAM, while introducing DAE further boosts performance by adaptively regulating the denoising strength. The highest IoU and Dice scores are achieved when combining all modules. Qualitative results in Fig. 10 further validate these findings: the reconstructed latent features become progressively clearer and more structured as additional modules are incorporated. Moreover, as shown in Tab. 6, even without fine-tuning the decoder, the proposed method improves performance on low-quality images by approximately 7 percentage points, while maintaining robustness on high-quality inputs. These results underscore the effectiveness and generalization capability of our generative enhancement framework.
@@ 5.3.2 Effect of Fine-tuning SAM
Figure 10: Qualitative visualization of the enhanced latent features generated by different channel expansion methods. Our proposed CRE method (c) produces more salient features.
@@ 5.3.3 Analysis of channel expansion methods
(a) Additional encoder and decoder 0.449 0.586 0.620
(b) New head and tail layers 0.617 0.705 0.783
@@
5.3.4 Analysis of DAE method
5.3.5 Analysis of LoRA ranks
We integrate learnable LoRA layers into the pre-trained denoising U-Net to enable efficient fine-tuning. To assess the influence of different LoRA ranks on segmentation performance, we conduct experiments with ranks of 4, 8, and 16, as reported in Tab. 9. The results show that a rank of 8 achieves the best overall performance while keeping the number of learnable parameters at a moderate level, demonstrating a favorable trade-off between accuracy and efficiency.
5.3.6 Analysis of hyperparameter $\gamma$
@@
Figure 13: Performance comparison of interactive segmentation with varying quantities of input points on the unseen ECSSD dataset. GleSAMs consistently outperform SAMs across a range of point counts, demonstrating a more significant improvement.
@@ 6.3 Comparison of Other Prompts
6.4 Analysis of computational requirements.
As shown in Tab. 11, we report detailed training and inference comparisons among our GleSAM++, RobustSAM, and SAM. GleSAM++ achieves a favorable balance between performance and efficiency compared with SAM and RobustSAM. Note that GleSAM++ brings negligible extra computation burden to SAM, with about 3% increase in model parameters and incurs a slight trade-off in inference speed (0.06 s). Based on SAM, the model can be efficiently optimized within 16 hours using four GPUs, indicating excellent training scalability.
@@ 7.2 Future and Prospects
The proposed GleSAM++ opens several promising directions for future exploration: (1) the generative latent-space enhancement paradigm can be further extended to other vision tasks beyond image segmentation, such as object detection, video object segmentation and tracking, as well as 3D vision tasks. (2) It can also be generalized to other foundation models (such as DINO series [DINO-representation, dinov2, dinov3]) to improve their robustness and adaptability under challenging conditions. (3) Our approach can further be explored to implemented into multimodal models, such as vision-language foundation models (VLM) [clip] or multi-modal large language models (MLLM) [qwen2, internvl], holding great potential for developing degradation-aware multimodal agents capable of both perception and reasoning under real-world degradations. We hope this work inspires further research on enhancing model robustness, ultimately contributing to the reliable deployment of vision systems in real-world scenarios.
@@ 7.3 Limitations
Despite the strong experimental performance of GleSAM++, our method still has several limitations that warrant further investigation. First, the generative latent-space enhancement paradigm carries an inherent risk of feature hallucination. In cases of extreme degradation where structural information is irretrievably lost, the latent diffusion model (LDM) may synthesize visually plausible but semantically inconsistent representations to compensate for missing details. Such behavior can result in confident yet incorrect segmentations, which pose potential risks in safety scenarios that demand high-fidelity reconstruction. Second, while the used one-step generative enhancement significantly mitigates the computational overhead typical of diffusion-based frameworks (with only 0.06 s additional latency over SAM), it still leaves room for further optimization. Future research could explore more lightweight generative mechanisms or adaptive inference strategies to reduce latency and memory consumption without sacrificing reconstruction fidelity or segmentation accuracy.

### v2

@@ 5.3 Ablation study
As summarized in Tab. 5, we conduct an ablation study to assess the individual contributions of the proposed modules in GleSAM++, namely the Feature Distribution Alignment (FDA), Channel Replication and Expansion (CRE), and Degradation-aware Adaptive Enhancement (DAE). Here, “Gle” denotes the base framework equipped with generative latent space enhancement. To ensure a fair comparison of reconstruction capability, all experiments are performed by fine-tuning only the U-Net backbone.
@@ 5.3.1 Effect of each component
The quantitative results demonstrate that each additional module consistently improves segmentation accuracy across all degradation levels and datasets. Notably, integrating FDA and CRE significantly bridges the gap between the pre-trained latent diffusion model and SAM, while introducing DAE further boosts performance by adaptively regulating the denoising strength. The highest IoU and Dice scores are achieved when combining all modules. Qualitative results in Fig. 10 further validate these findings: the reconstructed latent features become progressively clearer and more structured as additional modules are incorporated. Moreover, as shown in Tab. 6, even without fine-tuning the decoder, the proposed method improves performance on low-quality images by approximately 7 percentage points, while maintaining robustness on high-quality inputs. These results underscore the effectiveness and generalization capability of our generative enhancement framework.
@@ 5.3.2 Effect of fine-tuning SAM
Figure 11: Qualitative visualization of the enhanced latent features generated by different channel expansion methods. Our CRE method (c) produces more salient features.
@@ 5.3.3 Analysis of channel expansion methods
(a) Learnable projection layers 0.449 0.586 0.620
(b) Training head and tail layers from scrith 0.617 0.705 0.783
@@
5.3.4 Analysis of DAE method
Figure 12: Sensitivity analysis of the DAE mechanism under different perturbation magnitudes on the predicted degradation score.
To further evaluate the robustness of DAE under imperfect score estimation, we perform an additional sensitivity analysis by perturbing the predicted degradation score with different magnitudes. As shown in Fig. 12, the interpolation-based mapping exhibits only a mild and gradual performance drop as the perturbation increases, whereas directly using $s$ results in a much steeper degradation trend. This observation indicates that the proposed mapping strategy effectively regularizes the influence of score errors and prevents them from being excessively amplified during denoising. Therefore, beyond improving the average reconstruction quality, the proposed DAE design also provides stronger robustness and bounded sensitivity to inaccurate degradation prediction.
5.3.5 Analysis of denoising steps
We further conduct a systematic ablation study by varying the number of denoising steps from 1 to 5 under three degradation levels. We claim that one-step denoising provides the most favorable balance between segmentation performance and computational efficiency. As reported in Tab. 10, the best segmentation performance is achieved within a very small number of denoising steps. Importantly, the performance does not continue to improve as the number of steps increases. Instead, it gradually declines after the third step. This trend suggests that a limited amount of iterative refinement is beneficial for restoring degraded semantic features, whereas excessive denoising tends to cause over-correction or over-smoothing, thereby weakening boundary details and discriminative cues that are critical for accurate segmentation.
We also report the inference efficiency under different denoising steps. Although increasing the denoising process from 1 step to 3 steps brings marginal performance gains, the additional computational cost and latency are disproportionately larger. Therefore, 1-step denoising offers the most favorable trade-off between segmentation performance and inference efficiency. This observation further supports our design choice of single-step denoising as a practical and efficient default setting.
Table 10: Ablation study on the number of denoising steps under three degradation levels. Single-step denoising provides the best accuracy-efficiency trade-off.
Step LQ3 LQ2 LQ1 Average Latency
IoU Dice IoU Dice IoU Dice IoU Dice (s)
Baseline 0.525 0.643 0.608 0.715 0.676 0.770 0.603 0.709 0.32
1 0.730 0.824 0.773 0.855 0.811 0.882 0.771 0.854 0.38↑0.06
2 0.747 0.836 0.788 0.865 0.822 0.889 0.786 0.863 0.51↑0.19
3 0.748 0.836 0.790 0.867 0.821 0.888 0.786 0.864 0.56↑0.24
4 0.739 0.829 0.783 0.861 0.813 0.883 0.778 0.858 0.66↑0.34
5 0.729 0.821 0.770 0.853 0.802 0.876 0.767 0.850 0.81↑0.49
@@ 5.3.5 Analysis of denoising steps
5.3.6 Analysis of LoRA ranks and hyperparameter $\gamma$
In this part, we analyzed in detail the effects of two hyperparameters: LoRA ranks and $\gamma$. We integrate learnable LoRA layers into the pre-trained denoising U-Net to enable efficient fine-tuning. To assess the influence of different LoRA ranks on segmentation performance, we conduct experiments with ranks of 4, 8, and 16, as reported in Tab. 9. The results show that a rank of 8 achieves the best overall performance while keeping the number of learnable parameters at a moderate level, demonstrating a favorable trade-off between accuracy and efficiency.
@@ 5.3.6 Analysis of LoRA ranks and hyperparameter $\gamma$
5.3.7 Analysis of two-stage learning
We further analyze the effect of the proposed two-stage learning strategy. Compared with single-stage joint optimization, this design reduces the interference between feature restoration and segmentation learning, leading to more stable and effective training. As shown in Tab. 11, we report the averaged results on the four datasets and find that two-stage learning consistently improves segmentation performance across all degradation levels. Importantly, even when the number of training iterations for single-stage training is substantially increased, its performance still remains inferior to that of the proposed two-stage strategy. This suggests that the benefit of the two-stage design does not simply arise from longer training, but from a more effective optimization process that better decouples feature restoration and segmentation learning.
Table 11: Effect of the two-stage learning strategy, which leads to more stable optimization and stronger robust segmentation performance.
Method Training LQ3 LQ2 LQ1
Time IoU Dice IoU Dice IoU Dice
Single-stage Training 11h 0.617 0.734 0.639 0.751 0.648 0.758
24h 0.635 0.749 0.657 0.765 0.664 0.770
Two-stage Training (Ours) 16h 0.658 0.765 0.700 0.797 0.719 0.812
@@ 5.3.7 Analysis of two-stage learning
Figure 15: Performance comparison of interactive segmentation with varying quantities of input points on the ECSSD dataset. GleSAMs consistently outperform SAMs across a range of point counts, demonstrating a more significant improvement.
@@ 6.3 Comparison of other prompts
Figure 16: Representative failure cases of GleSAM++ under severe degradations. The main errors fall into two categories: semantically incorrect but visually plausible reconstruction, and failure caused by extreme information loss.
6.4 Analysis of computational requirements
GleSAM++ achieves a favorable trade-off between robustness and efficiency compared with SAM and RobustSAM. Although our framework incorporates a pre-trained diffusion U-Net, the computational overhead remains limited because the generative backbone is used mainly as a source of pre-trained priors and is adapted in a parameter-efficient manner, rather than being fully fine-tuned. Moreover, our method performs enhancement directly in the latent space and adopts single-step denoising at inference time, which avoids both the heavy cost of pixel-space restoration and the iterative sampling overhead of conventional diffusion models. Consequently, as shown in Tab. 13, GleSAM++ requires only 47M learnable parameters and can be trained within 16 hours on 4 GPUs, while adding only 0.06 s latency over SAM during inference. These results demonstrate that the introduced generative module remains computationally practical and does not substantially increase the overall system complexity.
6.5 Faliure case analysis
Fig. 16 shows representative failure cases of GleSAM++ under severe degradations. We observe two major failure modes. The first is visually plausible but semantically incorrect reconstruction. In such cases, the enhanced latent representation appears structurally reasonable, but due to insufficient semantic guidance under heavy degradation, the restored features deviate from the true object semantics, resulting in inaccurate object shapes or incomplete segmentation regions. The second is failure caused by extreme information loss. When the degraded input loses too much structural and boundary information, the enhancement module can no longer reliably recover the correct target representation, leading to obvious reconstruction and segmentation errors.
@@ 7.2 Future and Prospects
The proposed GleSAM++ opens several promising directions for future exploration: (1) the generative latent-space enhancement paradigm can be further extended to other vision tasks beyond image segmentation, such as video object segmentation with SAM2 [sam2]. Although our framework is structurally compatible with SAM2, the current study focuses on robustness at the frame-level representation stage under degraded-image settings, rather than in a full video segmentation scenario. Specifically, the latent enhancement is applied before mask decoding and does not explicitly modify SAM2’s temporal memory propagation mechanism. Therefore, the current results mainly verify backbone compatibility, while the effect of latent enhancement on temporal consistency across consecutive degraded frames remains to be studied. A rigorous investigation of this issue would require a dedicated degraded-video benchmark and suitable temporal evaluation metrics. (2) It can also be generalized to other foundation models (such as DINO series [DINO-representation, dinov2, dinov3]) to improve their robustness and adaptability under challenging conditions. (3) Our approach can further be explored to implemented into multimodal models, such as vision-language foundation models (VLM) [clip] or multi-modal large language models (MLLM) [qwen2, internvl], holding great potential for developing degradation-aware multimodal agents capable of both perception and reasoning under real-world degradations. We hope this work inspires further research on enhancing model robustness, ultimately contributing to the reliable deployment of vision systems in real-world scenarios.
@@ 7.3 Limitations
Despite the strong experimental performance of GleSAM++, our method still has several limitations that warrant further investigation. First, the generative latent-space enhancement paradigm carries an inherent risk of feature hallucination. In cases of extreme degradation where structural information is irretrievably lost, the latent diffusion model (LDM) may synthesize visually plausible but semantically inconsistent representations to compensate for missing details (as analyzed in Sec. 6.5). Such behavior can result in confident yet incorrect segmentations, which pose potential risks in safety scenarios that demand high-fidelity reconstruction. Second, while the used one-step generative enhancement significantly mitigates the computational overhead typical of diffusion-based frameworks (with only 0.06 s additional latency over SAM), it still leaves room for further optimization. Future research could explore more lightweight generative mechanisms or adaptive inference strategies to reduce latency and memory consumption without sacrificing reconstruction fidelity.
@@ 8 Conclusion
Declarations
Conflict of interest statement: The authors have no relevant financial or non-financial interests to disclose.
Funding information: The authors did not receive support from any organization for the submitted work.
Data Availability: The LQSeg dataset constructed in this study is publicly available at HuggingFace: https://huggingface.co/guogq/GleSAM.
Compliance with Ethical Standards: This article does not contain any studies with human participants or animals performed by any of the authors.